OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (4, -3). y = 4 x = -3 x = 4 y = -3

OpenStudy (cwrw238):



OpenStudy (anonymous):

so its -3

OpenStudy (cwrw238):

yes y = -3

OpenStudy (anonymous):

thanks ho bout this.. Which of the following is the slope between the two points (0, -4) and (-2, 5)? -9/2 2/9 -2/9 9/2

OpenStudy (cwrw238):

slope = difference in y coordinates / difference in x coordinates = (5-(-4)) / (-2-0)

OpenStudy (anonymous):

so its A

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thankss do u mind helping with one more question

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

Choose the equation of the vertical line that passes through the point (-5, 9). y = -5 y = 9 x = -5 x = 9 i think its -5

OpenStudy (cwrw238):

yes x = -5

OpenStudy (anonymous):

thankss

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

i got one more question i want to make sure Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -1/2x + 5. y = -2x + 2 y = 2x - 14 y =1/2 x - 8 y = -1/2x - 4

OpenStudy (cwrw238):

an equation parallel to y = -1/2 x + 5 will gave the same slope the slope is the coefficient of x that is -1/2

OpenStudy (anonymous):

so its D

OpenStudy (cwrw238):

right